# Chain-of-Authorization: Internalizing Authorization into Large Language Models via Reasoning Trajectories

### 24 Mar 2026 v1

@@ Abstract
Large Language Models (LLMs) have become core cognitive components in modern artificial intelligence (AI) systems, combining internal knowledge with external context to perform complex tasks. However, LLMs typically treat all accessible data indiscriminately, lacking inherent awareness of knowledge ownership and access boundaries. This deficiency heightens risks of sensitive data leakage and adversarial manipulation, potentially enabling unauthorized system access and severe security crises. Existing protection strategies rely on rigid, uniform defense that prevent dynamic authorization. Structural isolation methods faces scalability bottlenecks, while prompt guidance methods struggle with fine-grained permissions distinctions. Here, we propose the Chain-of-Authorization (CoA) framework, a secure training and reasoning paradigm that internalizes authorization logic into LLMs’ core capabilities. Unlike passive external defneses, CoA restructures the model’s information flow: it embeds permission context at input and requires generating explicit authorization reasoning trajectory that includes resource review, identity resolution, and decision-making stages before final response. Through supervised fine-tuning on data covering various authorization status, CoA integrates policy execution with task responses, making authorization a causal prerequisite for substantive responses. Extensive evaluations show that CoA not only maintains comparable utility in authorized scenarios but also overcomes the cognitive confusion when permissions mismatches. It exhibits high rejection rates against various unauthorized and adversarial access. This mechanism leverages LLMs’ reasoning capability to perform dynamic authorization, using natural language understanding as a proactive security mechanism for deploying reliable LLMs in modern AI systems.
@@ 1 Main
Large Language Models (LLMs) have become the core cognitive component of modern artificial intelligence (AI) systems such as autonomous agents. By integrating internal parametric knowledge with external context from retrieved documents or tool-calling feedback, LLMs can perform complex downstream tasks such as reasoning [giadikiaroglou-etal-2024-reasoning, chen2025reasoning], question answering [kamalloo-etal-2023-evaluatingqa, yue2025qasurvey], and tool calling [li-2025-toolcall, masterman2024toolcall]. However, in processing information, LLMs often treat all accessible data equally, lacking the inherent ability to recognize knowledge ownership and access boundaries. This cognitive indiscriminateness hinders LLMs from establishing logical boundaries between data with different permissions, fundamentally limiting the reliability of AI systems based on them when processing sensitive information.
Figure 1: Overview of the Chain-of-Authorization (CoA) framework. The upper half illustrates the mechanistic design, in which permission labels are injected into the input, forcing the model to generate a structured authorization trajectory before the final response. The lower half depicts the learning paradigm, which fine-tunes the LLM with synthesized data reflecting three authorization states, ensuring the authorization policy is embedded within the LLM.
This lack of inherent cognitive ability exposes LLMs to severe security challenges, primarily manifested in the risks of unauthorized knowledge leakage and covert privilege escalation. Since LLMs cannot determine the data permissions during inference, they are highly susceptible to unintentionally exposing sensitive content to unauthorized users. Furthermore, when prompted by adversarial prompts, LLMs often violate security policies due to a lack of understanding of the required instruction permissions. Such permission violations not only cause model outputs to deviate from alignment principles but may also serve as a springboard for unauthorized access to interconnected systems, potentially triggering catastrophic data security crises.
Existing data protection strategies mainly rely on cleaning sensitive content during the pre-processing stage (such as de-identifying and rewriting personally identifiable information or copyrighted data), and introducing mechanisms such as differential privacy during the training [yudifferentially, mireshghallah2021privacyregularization] or inference stage [ginart2022submix, flemings2024pmixed], which aim to reduce the model’s memorization of sensitive information. However, these methods uniformly block the model’s access to sensitive information, compromising the integrity of knowledge and failing to support authorized access. This rigid defense mechanism is difficult to meet the needs of flexible access control. Especially within the Retrieval Augmentation Generation (RAG) framework [wutschitz2023rethinking], even when retrieved documents carry different permission labels, LLMs still lack a principled mechanism for dynamically evaluating and filtering authorized contextual information based on user permissions.
To address the aforementioned challenges, current attempts mainly fall into two categories (shown in Figure 1): structured isolation [tiwari2024ifc, segal2025domba, jayaraman2025permlm] and prompt guidance methods [liu2025sudolm, saha2025sudollm]. The former attempts to physically block unauthorized information flow by sharding data and training independent sub-models. However, this paradigm faces scalability bottlenecks that can lead to high computational and storage costs when handling complex permissions. In contrast, prompt-guidance methods use a single model and attempt to implement soft constraints via permissions embedded in system prompts. While these methods avoid repetitive training, they can be bypassed by adversarial inputs due to the decoupling of policy execution and internal inference [liu2023promptinjection, greshake2023promptinjection]. More importantly, they struggle to differentiate output content based on user permissions due to a lack of understanding of knowledge permissions.
This gap reveals a key question: Can access control policies be internalized into the reasoning process of LLMs, so that the model can proactively determine whether the internal and external knowledge it relies on is within the current user’s authorized scope before generating an answer?
@@
To address these challenges, we propose Chain-of-Authorization (CoA), a novel training and reasoning paradigm that internalizes authorization as an inherent cognitive capability of LLMs. Unlike existing methods that rely on external authorization, CoA systematically reconstructs the model’s information flow: on the input side, we structurally embed user permissions with knowledge and tool permission labels to establish a clear permission context; on the output side, we force the model to generate a structured authorization reasoning trajectory before generating the final answer, autonomously determining permission boundaries through resource review, identity resolution, and decision-making; during the training phase, we use structured data with multiple authorization states (permission match, permission mismatch, and no permissions) to fine-tune the model, integrate this ”permission-aware reasoning prefix” with downstream task logic through supervised fine-tuning. This paradigm transforms authorization decisions from passive filtering after generation into causal premises for generation, enabling the model to proactively define knowledge boundaries and perform compliant reasoning in dynamic and complex permission environments.
We conducted extensive evaluations on five datasets covering internal knowledge, external context, and tool calling, targeting models of varying sizes, including Qwen3-1.7B [qwen3technicalreport], Llama-3.1-8B-Instruct [grattafiori2024llama3herdmodels], and Mistral-7B-v0.3 [jiang2023mistral7b]. Results show that, compared to baseline methods, CoA maintains comparable accuracy in authorized scenarios and achieves a very high compliance rejection rate in unauthorized scenarios. CoA also demonstrates robustness against various adversarial attacks, effectively preventing the model from being misled into leaking unauthorized information. By visualizing the hidden-state representations on the WMDP [li2024wmdp] dataset, we revealed the intrinsic mechanism of CoA in the representation space. Finally, we empirically demonstrated the causal impact of internal reasoning trajectories on model security decisions by conducting targeted intervention experiments on key stages of CoA.
2 Internalizing Authorization via Reasoning Trajectories
@@
The integration of LLMs as the cognitive core of modern AI systems has fundamentally shifted how information is processed and managed. In these systems, the model serves as the central component, combining internal parametric knowledge with external contextual feedback to generate textual responses or tool-calling decisions. However, despite their advanced reasoning capabilities, these models often indiscriminately utilize all available information, lacking an intrinsic recognition of the access boundaries of the knowledge used or the decisions performed. To address this limitation, we re-examine the interaction between LLMs and the environment from the perspective of controlled information flow. We no longer view the operation of LLMs as a simple query-response sequence, but rather as a cognitive process of observing and controlling specific information flows within strict permission boundaries.
2.1 Formalizing Authorization and the Challenge of Permission Mismatch
@@
Typically, the interaction between LLMs as the cognitive core and the environment can be formalized as a conditional probability distribution over outputs given a specific input. The complete input sequence $X\in\mathcal{X}$ received by LLMs encompasses the knowledge needed to prompt the model to generate responses, including the user prompt $Q$, the external context $E$, and the available tools $G$, while the output sequence $Y\in\mathcal{Y}$ represents the natural language response or tool calling decision generated by the LLM. This process is controlled by the LLM’s parameter $\theta$, which can be represented as a conditional probability $\pi_{\theta}(Y|X)$. This mapping represents the fundamental information processing capability of LLMs, combining internal parametric knowledge with external information to respond to user prompts.
@@
To systematically implement authorization, we introduce an explicit set of permission tags $\mathcal{C}$ to mark the access boundaries of the information space. Specifically, a user query $Q$, the available tools $G$, and the external context $E$ are associated with permission labels $C_{q},C_{g},C_{e}\subseteq\mathcal{C}$, respectively. These labels define the requirements for accessing the corresponding information. Users are also assigned with permission labels $C_{u}\subseteq\mathcal{C}$ to indicate their authorization status. We abstract complex real-world access boundaries as a set-inclusion relationship, where authorization is determined by whether the user’s permission set fully covers the permissions required by the target task. We define a policy function $\phi:\mathcal{C}\times\mathcal{C}\rightarrow\{0,1\}$ to evaluate the information flow for managing this topology relationship, where $\phi(\{C_{q}\cup C_{g}\cup C_{e}\},C_{u})=1$ indicates that the user has the right to access all the information required to respond to the prompt $X$, otherwise $\phi(\{C_{q}\cup C_{g}\cup C_{e}\},C_{u})=0$.
@@
where $\delta_{rej}$ denotes a degeneration distribution concentrated on refusal responses (i.e., ”Sorry, I cannot tell you this.”). To generate a response that respects access boundaries, the service must base the authorization decision on both the user’s permissions and the knowledge implied by the query.
Cognitive challenges user complex permissions: Although the final decision of the policy function is binary (compliant output or denial blocking), the underlying reasons for triggering security constraints ($\phi=0$) and the cognitive challenges they pose to LLMs are highly heterogeneous. Specifically, the denial state encompasses two distinct distributions in the feature space: one is complete lack of permissions, where the user has no internal permission labels, in which case $C_{u}=\{c_{pub}\}$ (a special public label), and the model only needs to establish a solid zero-trust baseline; the other is permission mismatch, where the user holds some internal permission labels, but these labels do not fully cover the permission required for the current prompt, in which case $C_{u}\nsubseteq(C_{q}\cup C_{e}\cup C_{g})$ and $C_{u}\neq\{c_{pub}\}$. The latter requires the model to have a fine-grained understanding of permission relationships to prevent overgeneralization of internal permissions.
@@
2.2 Internalizing Policy Function via Chain-of-Authorization
To address the limitations of external security mechanisms, we propose explicitly modeling the policy function $\phi$ as a structured reasoning process and internalizing it within the LLM’s generation process. Specifically, we model the language model $\pi_{\theta}(Y|X,C)$ and the policy function $\phi(\cdot,\cdot)$ together as a joint distribution $\pi_{\theta}(Y,T|X;C)$ with reasoning trajectories, where the trajectory $T$ explicitly carries the execution process of the policy function $\phi(\cdot,\cdot)$. Thus, we treat authorization as an indivisible cognitive step in the generation process, rather than as an external decision made before or after generation.
To overcome the limitations imposed by external components or shallow soft constraints, we systematically reconstruct the information flow of LLM through mechanism design and learning paradigms. This reconstruction aims to embed the policy function $\phi$ into LLMs’ reasoning process, transforming it from a passive external barrier into an endogenous cognitive ability. In terms of mechanism design, we model the language model $\pi_{\theta}(Y|X,C)$ and the policy function $\phi(\cdot,\cdot)$ together as a joint distribution $\pi_{\theta}(Y,T|X;C)$, where the reasoning trajectory $T$ explicitly carries the execution process of the policy function $\phi(\cdot,\cdot)$. In terms of the learning paradigm, we place authorization reasoning and task outputs within the same sequence prediction framework, enabling the model to naturally acquire the corresponding authorization decision logic as it learns to complete downstream tasks.
2.2.1 Mechanistic Design: Permission-Aware Input-Output Reformulation
To overcome the limitation of LLMs lacking explicit awareness of information ownership, we first establish a structured causal relationship between the LLM’s input and output spaces. At the input end, we explicitly inject the permission context into the LLM’s input sequence. Specifically, we first inject the user’s permission labels $C_{u}$ into the system prompt, thereby maintaining semantic isolation from user input and preventing direct manipulation by the user. Furthermore, the retrieved context and available tools are also assigned with corresponding permission labels $C_{e}$ and $C_{g}$ to explicitly identify their knowledge ownership or access level. The reconstructed complete input $X$ is formalized as
@@ 2.2.1 Mechanistic Design: Permission-Aware Input-Output Reformulation
where $C_{u}$ and $C_{g}$ represent the permission labels of the user and tools, respectively. They are dynamically injected into the system prompts as global constraints for inference. $Context(E,C_{e})$ then explicitly annotates the external knowledge with permission labels $C_{e}$. After this reconstruction, the input received by LLMs is no longer a simple instruction, but an authorization scenario with a clear identity and permission relationship structure.
<think>
The question is about [Prompt Permission] // Resource review stage
[Decision] // Decision-making stage
</think>
[Response] // Final output to user
However, establishing permission boundaries merely at the input end is insufficient to induce secure generation behavior. Therefore, at the output end, we introduce a Chain-of-Authorization (CoA) mechanism, establishing authorization as a necessary step before generating a substantial response $Y$. Based on the uniformly structured input $X$ described in Equation 1, we model this joint generation process as:
$\pi_{\theta}(Y,T|X)=\pi_{\theta}(T|X)\cdot\pi_{\theta}(Y|X,T)$ (2)
@@
where $T_{res}$ represents the resource review stage, used to analyze the permission of the knowledge ($C_{q}$), retrieved context ($C_{e}$), and tools ($C_{g}$) required to respond to the prompt. $T_{id}$ corresponds to the identity resolution stage, used to identify and interpret the current user’s permission status ($C_{u}$). $T_{dec}$ represents the final decision-making stage, providing a clear authorization conclusion based on the previous two stages and constraining subsequent responses.
An extremely simple illustration is shown in Figure 2, where the $2$ to $4$ lines demonstrate the resource review, identity resolution, and decision-making stage, respectively. Illustrations of external context authorization and tool-calling authorization are shown in Figure 8 (later in Section 5). Through this structured joint reconstruction, we establish an intrinsic security reasoning paradigm with strict temporal dependencies for the model.
2.2.2 Learning Paradigm: Internalizing Authorization via Supervised Fine-tuning
After establishing the structured representation, the new challenge becomes how to naturally internalize this authorization policy into the model parameters $\theta$. The model must learn not only to generate text, but also to judge whether a user has permission to perform a given task. To this end, we categorize the relationship between the user permissions $C_{u}$ and task requirements $C_{req}=C_{q}\cup C_{e}\cup C_{g}$ into three authorization states, and construct corresponding instruction fine-tuning data $(X,T,Y)$ from the downstream dataset $D=(C,X,Y)$ with permission requirements, thus forcing the model to learn precise permission boundaries:
  • Matched authorization ($C_{req}\subseteq C_{u}$): When the user’s permission fully encompasses the task’s permission requirements, the model is trained to generate a valid authorization trajectory $T$ and execute the downstream task, thereby preserving system utility.
  • Mismatched authorization ($C_{req}\nsubseteq C_{u}$ and $C_{u}\neq\{c_{pub}\}$): In scenarios where a user possesses partial permissions but requests restricted information, CoA forces the model to explicitly evaluate label discrepancies during the decision stage ($T_{dec}$). By accurately identifying these unauthorized attempts and redirecting the output $Y$ to the rejection distribution $\delta_{rej}$, this mechanism overcomes the cognitive confusion inherent in prompt-guidance methods and provides the model with robustness against unauthorized access.

### 03 Apr 2026 v2

# Chain-of-Authorization: Embedding authorization into large language models

@@ Abstract
Although Large Language Models (LLMs) have evolved from text generators into the cognitive core of modern AI systems, their inherent lack of authorization awareness exposes these systems to catastrophic risks, ranging from unintentional data leakage to unauthorized command execution. Existing defense mechanisms are fundamentally decoupled from internal reasoning, rendering them insufficient for the complex security demands of dynamic AI systems. Here, we propose the Chain-of-Authorization (CoA) framework, a paradigm that internalizes access control as a foundational cognitive capability. By systematically redesigning the input-output format and fine-tuning the model on synthesized data with complex permission topologies, CoA forces the model to generate a structured authorization trajectory as a causal prerequisite for any substantive response or action, thereby enabling LLMs to internalize access boundaries within dynamic reasoning environments. CoA maintains high utility in authorized scenarios while achieving high rejection rates of unauthorized prompts and robust defense against diverse adversarial attacks. By embedding authorization directly into the reasoning process, CoA provides a principled architectural blueprint for deploying secure LLMs as the cognitive cores of modern AI systems.
@@ 1 Main
The integration of Large Language Models (LLMs) as the cognitive core of modern artificial intelligence (AI) systems has fundamentally shifted the paradigm of human-computer interaction, driving applications ranging from conversational chatbots to high-level agents. This paradigm shift enables LLMs to perform an increasingly complex array of tasks, ranging from text generation to interacting directly with private desktops, executing system-level commands, and managing sensitive enterprise data [Wang2024agentsurvey, li2024personalagents, luo2025LLMagent]. As these systems move from “speaking” to “acting”, the security of their internal decision-making becomes a matter of systemic stability. Navigating such a diverse spectrum of tasks with varying security requirements inherently demands strict privilege separation to isolate risks. However, a critical architectural flaw remains: as the “central processor” of the system, LLMs do not inherently enforce privilege separation, leaving them blind to the access boundaries governing the information they process and the tools they wield.
Figure 1: Overview of the Chain-of-Authorization (CoA) framework. The upper half illustrates the mechanistic design, and the lower half illustrates the learning paradigm.
This authorization blindness hinders LLMs from establishing logical firewalls between data with varying permission levels. For example, in the context of autonomous agents, this manifests as the “Confused Deputy” problem [ji2026taming, siu2026framework, liu2026visual], where agents are misled by malicious instructions embedded in a public webpage or document, inadvertently seize its high-privilege system access to execute a low-privilege, malicious request [greshake2023promptinjection, liu2023promptinjection]. This lack of understanding of authorization not only heightens the risk of sensitive data leakage but also turns the LLM into a springboard for unauthorized access to interconnected systems, triggering severe security crises.
Current data protection strategies are largely passive and external. They rely on pre-processing (de-identifying data) or post-inference filtering, which aims to reduce the model’s memorization of sensitive information [yudifferentially, mireshghallah2021privacyregularization, ginart2022submix, flemings2024pmixed]. However, these rigid defense mechanisms are insufficient for the dynamic, multi-tenant environments in which AI systems now operate. Besides, prompt guidance methods [liu2025sudolm, saha2025sudollm, almheiri2025role] are fundamentally decoupled from the model’s internal reasoning; they can be bypassed by carefully designed adversarial prompts or “jailbreaks” that manipulate the model’s semantic understanding to circumvent shallow constraints [chao2025pair, zeng-etal-2024pap, liu2024autodan, zhou2024easyjailbreak]. Furthermore, structured isolation, which trains independent submodels on sharded data [tiwari2024ifc, jayaraman2025permlm, segal2025domba], faces severe scalability bottlenecks as permission structures grow more complex. The central challenge remains: can access control policies be internalized into LLMs’ reasoning processes, so that the model proactively determines its own authorized scope before generating an answer or executing an action?
@@
To address this, we propose Chain-of-Authorization (CoA), a novel training and reasoning paradigm that transforms authorization from an external filter into an inherent cognitive capability. Unlike previous methods, CoA systematically reconstructs the models’ information flow. On the input side, we structurally embed permission context, including user credentials, tool access levels, and context-specific labels, into the system prompt to establish a clear permission landscape. On the output side, we force the model to generate an explicit, structured authorization reasoning trajectory before arriving at the final response. This trajectory requires the model to perform a resource review, resolve the user’s identity, and make a formal decision, thereby making authorization a causal premise for generation.
We conducted extensive evaluations across five diverse datasets covering internal parametric knowledge, external retrieval-augmented context, and tool-calling scenarios. We tested CoA on various model architectures, including Qwen3-1.7B [qwen3technicalreport], Llama-3.1-8B [grattafiori2024llama3herdmodels], and Mistral-7B [jiang2023mistral7b], demonstrating that our approach maintains comparable utility in authorized stats while achieving a near-zero success rate for unauthorized access attempts. Through visualization of the hidden layer representations, we reveal the intrinsic mechanism of CoA, showing how it separates authorized and unauthorized semantic regions within the model’s representation space. Finally, we empirically demonstrate the causal impact of these reasoning trajectories, proving that CoA enables next-generation AI agents to function as secure, self-auditing entities within highly sensitive interconnected environments.
2 Authorization via Reasoning Trajectories
@@
The integration of LLMs as the cognitive core of modern AI systems has fundamentally shifted how information is processed and managed. In these systems, the model serves as the central component, combining internal parametric knowledge with external contextual feedback to generate textual responses or tool-calling decisions. However, despite their advanced reasoning capabilities, these models often indiscriminately utilize all available information, lacking an intrinsic recognition of the access boundaries of the knowledge used or the decisions performed. In this section, we will describe the theoretical architecture and implementation path of the Chain-of-Authorization framework. We first formally define the data access control problem in the context of Large Language Models (LLM), and then delve into the core mechanism design and the corresponding learning paradigm.
2.1 Formalizing Authorization for LLMs
@@
Typically, the interaction between LLMs as the cognitive core and the environment can be formalized as a conditional probability distribution over outputs given a specific input. The complete input sequence $X\in\mathcal{X}$ received by LLMs encompasses the knowledge needed to prompt the model to generate a response, including the user prompt $Q$, the external context $E$, and the available tools $G$, while the output sequence $Y\in\mathcal{Y}$ represents the natural language response or tool calling decision generated by the LLM. This process is controlled by the LLM’s parameter $\theta$, which can be represented as a conditional probability $\pi_{\theta}(Y|X)$. This mapping represents the fundamental information-processing capability of an LLM, combining internal knowledge parameters with external information to respond to user prompts.
@@
To systematically implement authorization, we introduce an explicit set of permission labels $\mathcal{C}$ to mark the access boundaries of the information space. Specifically, a user query $Q$, the available tools $G$, and the external context $E$ are associated with permission labels $C_{q},C_{g},C_{e}\subseteq\mathcal{C}$, respectively. These labels define the requirements for accessing the corresponding information. Users are also assigned with permission labels $C_{u}\subseteq\mathcal{C}$ to indicate their authorization status. Based on this, we abstract the complex access boundaries in reality as a set of matching relationships between user permissions and the permissions required by the target task. We define a policy function $\phi:\mathcal{C}\times\mathcal{C}\rightarrow\{0,1\}$ to evaluate the information flow for managing this topology relationship, where $\phi(\{C_{q},C_{e}\},C_{u})=1$ indicates that the user has the right to access all the information required to respond to the prompt $X$, otherwise $\phi(\{C_{q},C_{e}\},C_{u})=0$.
@@
where $\delta_{rej}$ denotes a degeneration distribution concentrated on refusal responses (i.e., “Sorry, I cannot tell you this.”). To generate a response that respects access boundaries, the service must base the authorization decision on both the user’s permissions and the knowledge implied by the query.
Cognitive challenges under complex permissions: Although the final decision of the policy function is binary (compliant output or denial blocking), the underlying reasons for triggering security constraints ($\phi=0$) and the cognitive challenges they pose to LLMs are highly heterogeneous. Specifically, the denial state encompasses two distinct distributions in the feature space: one is complete lack of permissions, where the user has no internal permission labels, in which case $C_{u}=\{c_{pub}\}$ (a special public label), and the model only needs to establish a solid zero-trust baseline; the other is permission mismatch, where the user holds some internal permission labels, but these labels do not fully cover the permission required for the current prompt, in which case $C_{u}\nsubseteq(C_{q}\cup C_{e}\cup C_{g})$ and $C_{u}\neq\{c_{pub}\}$. The latter requires the model to have a fine-grained understanding of permission relationships to prevent valid but wrong permissions.
@@
2.2 Chain-of-Authorization Framework
As shown in Figure 1, to overcome the limitations imposed by external components or shallow soft constraints, we systematically reconstruct the information flow of LLM through mechanism design and learning paradigms. This reconstruction aims to embed the policy function $\phi$ into LLMs’ reasoning process, transforming it from a passive external barrier into an endogenous cognitive ability. In terms of mechanics design, we explicitly embed permission labels into the input and enforce a structured reasoning process to evaluate these constraints prior to response generation. In terms of the learning paradigm, we synthesize data with different permission topologies, and place authorization reasoning and task outputs within the same sequence prediction framework, enabling the model to naturally acquire the corresponding authorization decision logic as it learns to complete downstream tasks.
2.2.1 Input-Output Reformulation
To overcome the limitation of LLMs lacking explicit awareness of information ownership, we first establish a structured causal relationship between the LLM’s input and output spaces. At the input end, we explicitly inject the permission context into the LLM’s input sequence. Specifically, we first inject the user’s permission labels $C_{u}$ into the system prompt, thereby maintaining semantic isolation from user input and preventing direct manipulation by the user. Furthermore, the retrieved context and available tools are also assigned with corresponding permission labels $C_{e}$ and $C_{g}$ to explicitly identify their knowledge ownership or access level. The reconstructed complete input $X$ is formalized as:
@@ 2.2.1 Input-Output Reformulation
where $C_{u}$ and $C_{g}$ represent the permission labels of the user and tools, respectively. They are dynamically injected into the system prompts as global constraints for inference. $Context(E,C_{e})$ then explicitly annotates the external knowledge with permission labels $C_{e}$, $\oplus$ denotes the concatenation operation. After this reconstruction, the input received by LLMs is no longer a simple instruction, but an authorization scenario with a clear identity and permission relationship structure.
However, establishing permission boundaries merely at the input end is insufficient to induce secure generation behavior. Therefore, at the output end, we introduce a Chain-of-Authorization (CoA) mechanism, establishing authorization as a necessary step before generating a substantial response $Y$. Based on the uniformly structured input $X$ described in Equation 1, we model this joint generation process as
$\pi_{\theta}(Y,T|X)=\pi_{\theta}(T|X)\cdot\pi_{\theta}(Y|X,T).$ (2)
@@
where $T_{res}$ represents the resource review stage, used to analyze the permission of the knowledge ($C_{q}$), retrieved context ($C_{e}$), and tools ($C_{g}$) required to respond to the prompt. $T_{id}$ corresponds to the identity resolution stage, used to identify and interpret the current user’s permission status ($C_{u}$). $T_{dec}$ represents the final decision-making stage, providing a clear authorization conclusion based on the previous two stages and constraining subsequent responses. Through this structured joint reconstruction, we establish an intrinsic security reasoning paradigm with strict dependencies for the model.
2.2.2 Data Synthesis and Fine-tuning
After establishing the structured representation, the next challenge is how to naturally integrate this authorization policy into the model. Based on whether the user has the permission to constrained data, we categorize the relationship between user permissions $C_{u}$ and task requirements $C_{req}=C_{q}\cup C_{e}\cup C_{g}$ into three typical authorization states, and construct corresponding instruction fine-tuning data $(X,T,Y)$ from the downstream dataset $D=(C,X,Y)$ with permission requirements, thus forcing the model to learn precise permission boundaries:
  • Matched authorization ($C_{req}\subseteq C_{u}$): When the user’s credentials fully encompass the task’s permission requirements, the model is trained to generate a valid authorization trajectory $T$ and execute the downstream task, thereby preserving system utility.
  • Mismatched authorization ($C_{req}\nsubseteq C_{u}$ and $C_{u}\neq\{c_{pub}\}$): In scenarios where a user possesses partial permissions but requests restricted information, CoA forces the model to explicitly evaluate label discrepancies during the decision stage ($T_{dec}$). By accurately identifying these unauthorized attempts and redirecting the output $Y$ to the rejection distribution $\delta_{rej}$, this mechanism overcomes the cognitive confusion inherent in prompt-guided methods and provides the model with robustness against unauthorized access.